OpenStudy (anonymous):

The sum of 3 consecutive EVEN integers is 174. What is the first integer?

OpenStudy (cwrw238):

if first integer is n we have n + n +2 + n + 4 = 174 solve for n

OpenStudy (anonymous):

I am confused on the solving part

OpenStudy (cwrw238):

on left hand side add the3 'n's and add 2 and 4 what do you get?

OpenStudy (cwrw238):

what is n + n + n?

OpenStudy (anonymous):

3n

OpenStudy (cwrw238):

yep and so LHS = 3n + 6 so we have 3n + 6 = 174 can you continue?

OpenStudy (anonymous):

subtract 6 from 174?

OpenStudy (cwrw238):

well almost - the equation must balance so you subtract 6 from both sides of the = 3n + 6 - 6 = 174 - 6 3n = 168 n = ?

OpenStudy (anonymous):

n=56?

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

thank you :)